OpenStudy (anonymous):

y varies directly as the square of x. When x = 5, y = 5. Find y when x = 10.

OpenStudy (anonymous):

x^2 is directly proportional to y so taking "k" as the proportionality constant we get x^2 = yk 5^2=5k 25=5k k=5 so when x=10 10^2=yk 100 = 5y therefore y=20...